OpenStudy (lexilove0828):

What is 5.8×1065.8×106 seconds in minutes? 3.48×10^4 ​​minutes 9.67×10^8​​ minutes 9.67×10^4 minutes 3.48×10^8 ​​minutes

OpenStudy (sooobored):

im going to assume that its \[5.8\times 10^6\]

OpenStudy (sooobored):

and im sure you know that there are 60 seconds in 1 minute

OpenStudy (lexilove0828):

thats not on there...

OpenStudy (sooobored):

5.8×1065.8×106 from your question

OpenStudy (sooobored):

if we want to make the conversion from seconds to minutes we can write it something like this \[\frac{5.8\times10^6 seconds}{}\frac{1 minute}{60 seconds}= \]

OpenStudy (sooobored):

hence the resulting solution is just \[5.8\times 10^6 \div 60\]